OpenStudy (anonymous):

Two end points of a line segment are (2, 8) and (-4, -8). What are the coordinates of the point on the line through which its bisector passes? (-2, 0) (-1, 0) (3, 8) (5, -6)

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Find the midpoint x-coordinate: (x1+x2)/2 = (2+(-4))/2 = -1 y-coordinate: (y1+y2)/2 = (8+(-8))/2 = 0 So the midpoint is (-1,0) Therefore, the bisector passes through (-1,0) So the answer is choice B